OpenStudy (anonymous):

An event lasts 5 hours and 30 minutes. How long did the event last in seconds?

OpenStudy (anonymous):

5*3600 + 30*60

OpenStudy (anonymous):

5400 seconds

OpenStudy (anonymous):

thanks you all ^.^

OpenStudy (anonymous):

1hr=60 minutes 5*60+30=330min 1min=60 sec calculate 330*60

OpenStudy (anonymous):

yw ^_^

OpenStudy (anonymous):

\[5*3600+30*60=19800 \text{ seconds}\]

OpenStudy (anonymous):

\[330*60=19800 \]

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (anonymous):

sry I got the wrong answer I guess I calculated wrong. >v< @sashiaaaa

OpenStudy (anonymous):

its ok i got it Christineroxx i got i it thank you anways

OpenStudy (anonymous):

kk and yw :p